OpenStudy (anonymous):

Hi, "Write a function that takes in two numbers and recursively multiplies them together.".. What does it mean ? if a = 3 and b = 2, does it mean : 3x2+2x1+1x0 ? or 3!x2! ? Thank you

OpenStudy (microbot):

maybe you are supposed to white a function that reproduces the * .means that u cant use multiply but + so if u want to multiply 2*3 ull have to do 2+2+2....maybe.... but not sure really

OpenStudy (anonymous):

Hi, thanks to stop by. ;) This comes from : 6.189 IAP 2011: Optional Recursion Exercises but they don't seem to say anything about operator use ?..

OpenStudy (microbot):

hmm never done or seen that course...so cant really help u...its just in many courses they show u or ask you to make that kind of an exercise , that's why i mentioned it best of luck with it

OpenStudy (anonymous):

What I think s meant is that if you have a multiplication like 4 * 2, you multiply it as follows: 4 * 2 = 2 + 3 * 2 3 * 2 = 2 + 2 * 2 2 * 2 = 2 + 1 * 2 1 * 2 = 2 This can be programmed with a recursive function. (Of course it'd be better to decrement the two instead of the four...)

OpenStudy (microbot):

well ye that's what i meant ....gj explaining it

OpenStudy (anonymous):

Yes I guess thats it. Thank you slotema ! ;)